# Patrick Cheridito and Moritz Weiss

### v1

@@ 5.2 Tactical traders
where $c\in\operatorname{\mathbb{R}}^{+}$ is a damping factor. The weighted volume imbalance is then defined by
@@
  • For $d^{M}\in\operatorname{\mathbb{R}}^{+},$ market buy orders arrive with intensity $d^{M}I^{+}$ and market sell orders arrive with intensity $d^{M}I^{-}.$
  • For $k\in\{1,2,\dots,D\},$ limit buy orders arrive with intensity $d^{k,L}I^{+}$ at $k$ ticks below the best ask price, and limit sell orders with intensity $d^{k,L}I^{-}$ at $k$ ticks above the best bid price.
  • If the spread is $j$ ticks, then for $k\in\{j,\dots,D\},$ cancellation of limit buy orders arrive with intensity $d^{k,C}I^{-}v^{b,{k-j+1}},$ at $k$ ticks below the best ask price, and and cancellations of limit sell orders arrive with intensity $d^{C,k}I^{+}v^{a,{k-j+1}},$ at $k$ ticks above the best bid price.
For $k\in\{1,\dots,D\},$ market orders, limit orders and cancellations arrive with sizes $1+\sigma^{M}|Z|,1+\sigma^{L,k}L|Z|$ and $1+\sigma^{C,k}|Z|$ respectively, where $Z$ is a random variable with a standard normal distribution for both noise and tactical traders. Order sizes are always rounded to the closest integer. Furthermore, both types of traders always cancel orders with the highest queue positions first, and no more orders can be canceled than exist at a given level.
5.3 Strategic traders
The strategic trader buys or sells a position in the market using limit and market orders.
  • The direction, either buy or sell, is drawn randomly at the start time $-\Delta t$ with probability $1/2$ each.
  • The trader sends a market order of size $\nu^{M}\in\operatorname{\mathbb{N}}$ at time steps $-\Delta t,-\Delta t+\Delta t^{M},-\Delta t+2\Delta t^{M},...$ and stops trading at terminal time $T.$
  • The trader sends a limit order of size $\nu^{L}\in\operatorname{\mathbb{N}}$ at time steps $-\Delta t,-\Delta t+\Delta t^{L},-\Delta t+2\Delta t^{L},...$ and stops trading at terminal time $T.$
@@ 6 Numerical results
In this section, we test our RL-algorithm based on the logistic-normal distribution (LN) in different simulated market environments and compare it with the two heuristic benchmarks from Section 3.4 and another RL-algorithm based on the Dirichlet distribution (DR), explained in Appendix C below. Our source code is available at https://github.com/moritzweiss/rlte.
@@ 6.1 Simulation setup
We simulate three increasingly complex market environments. The first market contains only noise traders and serves as a baseline. Price movements are entirely random, with no response from market participants. The second market includes both noise traders and tactical traders. In this setting, order flows react to order book imbalances, a behavior commonly observed in empirical studies of limit order book dynamics; see e.g. Stoikov (2018). The third market adds a strategic trader to the previous participants. It simulates a market that exhibits upward or downward price drift due to the presence of a strategic trader. This drift requires the execution algorithm to adapt its execution strategy by front-loading or back-loading orders. While such pronounced trends are rare in real markets, this environment tests whether the algorithm can detect directional signals from order book features and adjust its behavior accordingly. Although these simulations are stylized, they capture key characteristics of real-world market behavior. Importantly, our simulations differ significantly from historical market data replay because in our simulations, trades generate market impact. First, large market orders have direct impact since they consume the best levels on the opposite side of the order book. Second, limit orders have indirect impact as the tactical traders react to order book imbalance. Those effects cannot be incorporated in a simulation relying on historical data replay, as the data is fixed and does not react to the algorithm’s trading decisions. We compare both algorithms with the heuristic algorithms [^5] We refer to the LN and DR algorithms and the benchmark algorithms as execution algorithms. defined in Section 3.4, in each of the three market environments, and for small and large positions to be liquidated.
As mentioned in Section 3, the execution algorithms trade at discrete time steps $t_{n}=n\Delta t$ for $n\in\{0,1,\dots,N-1\}.$ For our simulations, we choose $\Delta t=15s$ for a total of $N=10$ steps, such that the total execution duration is $150s.$ We use a short simulation horizon to reduce computational overhead during training. This choice does not compromise the validity of our results, as the underlying logic and outcomes would also hold for longer execution horizons. In each market, the execution algorithms trade a small (20) and a large (60) number of lots. What can be considered a small or large number of lots depends on the average traded volume in the interval $[0,T].$ In our examples, the average traded volume over the time interval is 100 lots in each market. When adding a new trader to the market, we adjust the trading intensities of the other traders so that the overall average trading volume remains the same at 100 lots. Hence, 20 lots correspond to 20% and 60 lots correspond to 60% of the total traded volume. In all three markets, we start the simulation at time $-\Delta t,$ where at initialization, we set the order book to its long-term average shape. Therefore, when the execution algorithms start trading at $t=0,$ the order book has moved from its initial state to a random state. The average shape of the order book can be estimated with a Monte Carlo simulation, which is described in more detail in Appendix A.2. The specific parameter choices for each trading agent are discussed in Appendix A.1.
We use the two heuristic strategies of Section 3.4 as benchmarks and introduce an additional benchmark based on the Dirichlet distribution, which is explained in more Detail in Section C. It would also be feasible to compare against other trade execution benchmark algorithms, such as the volume-weighted-average price algorithm. However, our algorithm is not directly comparable with other reinforcement learning algorithms that have been mentioned in the introduction. This has two reasons. First, algorithms developed on simulations relying on historical data are not comparable because our market simulation is inherently different from historical market data replay. Second, none of the related research works use the same state and action spaces formulation. In particular, no other work views the order placement problem as a dynamic allocation problem.
@@
Figure 2: The black solid lines are the best bid and ask prices. The triangles indicate buy and sell market orders. The colors indicate volumes. Darker colors correspond to larger volumes. Buy limit orders have a blue color, and sell limit orders have a red color.
The test results are summarized in Table 1, which contains the expected value and the standard deviation of the sum of normalized rewards defined in (3) for all market environments, position sizes, and execution algorithms. For example, the reward of $-1.10$ for the SL algorithm at 60 lots in the market with noise traders means that on average, the algorithm sold each of the 60 lots at $-1.10$ ticks less than the initial bid price. The plots in Figure 3 show histograms of normalized rewards based on 10,000 simulated executions. The random seeds used for the processes of different trading agents are selected independently from those employed during training, thereby avoiding over-fitting to the training simulations.
6.2 Reinforcement learning setup
@@
We train the LN algorithm with Algorithm 1. As explained in Sections 3.13.2, the parameter $K$ controls how many order book levels the algorithm observes and how deep it may post orders into the order book. In our simulations, the price rarely moves more than five ticks throughout the execution horizon. Therefore, we choose $K=6$ for the dimension of the simplex $\operatorname{\mathbb{S}}_{K}.$ This means that the RL algorithm may place limit orders up to five ticks above the best bid price and observes the volumes on the first five levels on both sides of the book. To encode the policy $\pi_{\theta}$ and the value function $V_{\vartheta},$ we choose two separate feed-forward neural networks with two hidden layers, $\tanh$-activation functions in the hidden layers and no activation functions in the last layers. The output layer of the policy network has $K$ nodes, and the last layer of the value function network has one node. We start by setting the covariance matrix $\Sigma$ to the identity matrix and then progressively reduce its values according to the linear schedule (12). We choose an initial bias $b=(-1,-1,\dots,-1)\in\operatorname{\mathbb{R}}^{K}$ for the policy network. We obtain from (10) that
@@
$\operatorname{\mathbb{E}}\left[\log\left(\frac{a^{K}}{a^{k}}\right)\right]=1\quad\text{and}\quad\operatorname{\mathbb{E}}\left[\log\left(\frac{a^{k}}{a^{K}}\right)\right]=-1,\quad k=0,1,\dots,K-1,$
@@
showing that action $a_{K}$ is more likely than actions $a_{0},\dots,a_{K-1}.$ This means the algorithm does not place any orders at the start of the training cycle, which ensures that the algorithm observes complete trajectories up to the terminal time $T$ early in the training process, preventing it from getting stuck in local minima. Further implementation details concerning the training of the LN algorithm are discussed in Appendix B.
@@ 6.3 Logistic-normal vs Dirichlet distribution
First of all, we observe that in all three market simulations, the LN algorithm outperforms the DR algorithm, which is visible in Table 1 and Figure 3. Figure 4 shows the development of the episode returns over the full training cycle. The upper panel shows the results for 20 lots, and the lower panel shows the results for 60 lots. We observe that in all experiments, except when the agent executes 20 lots in the market only containing noise traders, the rewards for the LN algorithm converge faster than for the DR algorithm. This suggests that the LN algorithm is a superior alternative to the DR algorithm in this allocation problem. As the performance of the LN and DR algorithms is qualitatively similar, we focus our attention on the performance of the LN algorithm relative the the heuristic benchmarks in the following sections.
Table 1: The table shows the expected rewards and their standard deviations for the submit and leave algorithm (SL), the time-weighted average price algorithm (TWAP), the logistic-normal algorithm (LN) and the Dirichlet algorithm (DR) in three simulated markets and for small and large initial positions.
Market #Lots $\operatorname{\mathbb{E}}[\text{SL}]$ $\sigma[\text{SL}]$ $\operatorname{\mathbb{E}}[\text{TWAP}]$ $\sigma[\text{TWAP}]$ $\operatorname{\mathbb{E}}[\text{DR}]$ $\sigma[\text{DR}]$ $\operatorname{\mathbb{E}}[\text{LN}]$ $\sigma[\text{LN}]$
Noise 20 0.52 1.20 -0.05 0.94 0.58 0.90 0.65 0.89
60 -1.10 1.34 -1.40 0.97 -0.74 0.92 -0.69 0.91
Noise & Tactical 20 0.10 1.43 0.48 0.68 0.73 0.67 0.79 0.69
60 -3.36 0.99 -0.96 0.95 -0.41 0.63 -0.25 0.65
Noise & Tactical 20 -1.61 2.95 -0.31 3.03 0.86 1.92 1.16 2.08
& Strategic 60 -2.46 3.67 -1.40 3.46 -0.02 1.90 0.16 2.04
@@
Figure 3: The figures show reward densities of the submit and leave algorithm (SL), the time-weighted average price algorithm (TWAP), the logistic-normal algorithm (LN) and the Dirichlet algorithm (DR) for three simulated markets. The first row corresponds to 20 lots, and the second to 60 lots.
@@
Figure 4: Average episode return per batch during training for different environments and initial position sizes. The x-axis shows the number of gradient steps. The top row shows the returns for 20 lots. The bottom row shows the returns for 60 lots.
@@ 6.4 Market with noise traders
The first two rows of Table 1 demonstrate that the LN algorithm outperforms both heuristic benchmarks and the DR algorithm for 20 and 60 lots. This is also visible in the histograms in Figure 3(a). We remark that all algorithms in all markets have a high standard deviation. This is unavoidable as the algorithms operate in a highly stochastic environment. Sending a market order at the beginning of the execution period would result in a strategy with almost zero standard deviation; however, the execution price would be worse than the initial best bid and, therefore, worse than most of our execution strategies. For 20 lots, the LN algorithm’s histogram plot exhibits a bump at one and two ticks above the best bid price, indicating that it places limit orders deeper in the book and cuts losses when the price moves against it. For 60 lots, while the histograms of the LN and SL algorithms appear similar, the LN algorithm’s histogram exhibits less skew to the left, further suggesting that it effectively limits losses when the market moves away. The TWAP algorithm has the worst performance for both position sizes. Since the market does not react to volume imbalance, placing the entire position passively at the start has no downside.

### v2

@@ 5.2 Tactical traders
where $c\in\operatorname{\mathbb{R}}_{+}$ is a damping factor. The weighted volume imbalance is then defined by
@@
  • For $d^{M}\in\operatorname{\mathbb{R}}_{+},$ market buy orders arrive with intensity $d^{M}I^{+}$ and market sell orders arrive with intensity $d^{M}I^{-}.$
  • For $k\in\{1,2,\dots,D\},$ limit buy orders arrive $k$ ticks below the best ask price with intensity $d^{L,k}I^{+},$ and limit sell orders arrive $k$ ticks above the best bid price with intensity $d^{L,k}I^{-}.$
  • If the spread is $j$ ticks, then for $k\in\{j,\dots,D\},$ cancellation of limit buy orders arrive $k$ ticks below the best ask price with intensity $d^{C,k}I^{-}v^{b,{k-j+1}},$ and cancellations of limit sell orders arrive $k$ ticks above the best bid price with intensity $d^{C,k}I^{+}v^{a,{k-j+1}}.$
For $k\in\{1,\dots,D\},$ market orders, limit orders and cancellations of the noise traders arrive with sizes $1+\delta^{M}_{\text{noise}}|Z|,1+\delta^{L,k}_{\text{noise}}|Z|$ and $1+\delta^{C,k}_{\text{noise}}|Z|$ respectively, where $Z$ is a random variable with standard normal distribution for both noise and tactical traders. Similarly, orders of the tactical traders arrive with sizes $1+\delta^{M}_{\text{tactical}}|Z|,1+\delta^{L,k}_{\text{tactical}}|Z|$ and $1+\delta^{C,k}_{\text{tactical}}|Z|.$ Order sizes are always rounded to the closest integer. Both types of traders always cancel orders with the highest queue positions first. No more orders can be canceled than exist at any given price level.
5.3 Strategic trader
The strategic trader buys or sells a position in the market using market and limit orders.
  • The direction, either buy or sell, is drawn randomly at the start time $\Delta t$ with probability $1/2$ each.
  • The trader sends a market order of size $\nu^{M}\in\operatorname{\mathbb{N}}$ at time steps $\Delta t,-\Delta t+\Delta t^{M},-\Delta t+2\Delta t^{M},$... and stops trading at terminal time $T.$
  • The trader sends a limit order of size $\nu^{L}\in\operatorname{\mathbb{N}}$ at the best ask price at time steps $\Delta t,-\Delta t+\Delta t^{L},-\Delta t+2\Delta t^{L},$... and stops trading at terminal time $T.$
@@ 6 Numerical results
In this section, we test our RL algorithm based on the logistic-normal distribution (LN) in different simulated market environments and compare it with the two heuristic benchmarks from Section 3.4 and another RL algorithm based on the Dirichlet distribution (DR). Our source code is available at https://github.com/moritzweiss/rlte. Although our simulations are stylized, they capture key characteristics of real-world market behavior. Importantly, our simulations differ significantly from historical market data replay because in our simulations, trades generate market impact. First, large market orders have direct impact since they consume the best levels on the opposite side of the order book. Second, limit orders have indirect impact as the tactical traders react to order book imbalance. Those effects cannot be incorporated into a simulation that relies on historical data replay, because historical data cannot respond to the algorithm’s trading decisions.
@@ 6.1 Simulation setup
We simulate three increasingly complex market environments. The first market contains only noise traders who behave purely randomly and do not react to other market participants. The second market includes both noise traders and tactical traders. In this setting, order flows react to order book imbalances, a behavior commonly observed in empirical studies of limit order book dynamics; see e.g. Stoikov (2018). The third market adds a strategic trader to the previous participants. It simulates a market that exhibits upward or downward price drift due to the presence of a strategic trader. We will see that our RL algorithm learns to adapt its execution strategy by front-or back-loading orders.
As mentioned in Section 3, the execution algorithms trade at discrete time steps $t_{n}=n\Delta t$ for $n\in\{0,1,\dots,N-1\}.$ For our simulations, we choose $\Delta t=15s$ for a total of $N=10$ steps, such that the total execution duration is $150$ seconds. This is a typical time horizon in high-frequency trading. There is substantial market activity within 150 seconds. On average, roughly 1000 order book updates (limit orders, market orders, or cancellations) occur in that time horizon, and a volume of 100 lots is traded; see Appendix A.3. In Appendix C.3, we also evaluate the execution algorithms on a longer time horizon of 300 seconds, achieving good results. Training over a much longer time horizon is also possible, but would require more computational resources. Our simulations are relatively expensive, as documented in Appendix B.1. Our reinforcement learning algorithm can be viewed as a specialized order-placing algorithm that can be integrated into a long-term execution schedule, such as Almgren and Chriss (2001).
In each market, the execution algorithms trade a small (20) and a large (60) number of lots. What can be considered a small or large number of lots depends on the average traded volume in the interval $[0,T].$ In our examples, the average traded volume over the time interval is roughly 100 lots in the first two markets, and slightly more in the last one. Precise statistics on traded volumes are documented in Appendix A.3. Hence, 20 lots correspond to roughly 20% and 60 lots correspond to 60% of the total traded volume.
In all three markets, we start the simulation at time $\Delta t,$ where at initialization, we start the order book from its long-term average shape. Therefore, when the execution algorithms start trading at $t=0,$ the order book has moved from its initial state to a random state. The average shape of the order book can be estimated with a Monte Carlo simulation, which is described in more detail in Appendix A.2. The specific parameter choices for each trading agent are reported in Appendix A.1. We use the two heuristic algorithms of Section 3.4 as benchmarks and introduce an additional benchmark based on the Dirichlet distribution, described in Section B.2. One could also compare against other trade execution benchmark algorithms, such as the volume-weighted-average price algorithm. However, our algorithm is not directly comparable with other reinforcement learning algorithms from the literature since they do not use the same state and action spaces.
@@
Figure 2: The black solid lines are the best bid and ask prices. The triangles indicate buy and sell market orders. The size of the triangles corresponds to the size of the market orders. The colors indicate volumes. Darker colors correspond to larger volumes. Buy limit orders have a blue color, and sell limit orders have a red color.
6.2 Market impact study
Before evaluating the performance of the reinforcement learning trading algorithms, we first conduct a market-impact study to quantify how market and limit order placement influences price dynamics. In each experiment, we initialize the market simulation at time $t=-\Delta t$. At time $t=0$, we submit either a market sell order or a limit sell order, placed at the best ask price, with sizes of 10, 20, or 60 lots. Figure 3 reports the average evolution of the mid-price relative to its value at time zero, p(t)p(0),p(t)-p(0), up to the terminal time $T=150,$ for each market and for different position sizes. Averages are computed based on 10,000 market simulations.
For market orders, we observe a sharp price drop immediately after the order is placed, followed by a slight recovery across all market environments. Naturally, larger market orders lead to a larger price impact in all cases. The mid-price recovery after the price drop is stronger in the market with noise and tactical traders than in the market with only noise traders. This is because the imbalance reaction from tactical traders has a mean-reverting effect, which pushes the price back up after a price level in the order book is removed. For the market consisting of noise, tactical, and strategic traders, the prices do not recover as much as in the other cases because when the strategic traders are selling, they amplify the downward move by continuing to push prices lower. For the limit order study, it is noticeable that the 60 lot limit order has much more market impact in the market with tactical traders. This is because the large limit order on the first level makes the order book very imbalanced. Since tactical traders react to order book imbalance, they send more market sell and limit sell orders, which drives the price down. As for market orders, we observe that larger limit orders have a larger impact in all cases. We conclude that both market and limit orders have substantial market impact in our simulations. This differentiates our work from works that rely on historical data simulations and usually ignore market impact.
(a) Noise Traders
(b) Noise & Tactical Traders
(c) Noise, Tactical, & Strategic Traders
Figure 3: Mid price evolutions after a market or limit order of size 10, 20, or 60 lots is placed at time $t=0.$ The upper panel shows mid price evolutions after a market order was placed. The lower panel shows mid price evolutions after a limit order was placed. The legend indicates the order sizes in lots. Deeper color tones correspond to larger order sizes.
6.3 Reinforcement learning setup and evaluation
@@
We train the LN algorithm with Algorithm 1. As explained in Sections 3.13.2, the parameter $K$ controls how many order book levels the algorithm observes and how deep it may post orders into the order book. In our simulations, the price rarely moves more than five ticks during the execution period. Therefore, we choose $K=6$ for the dimension of the simplex $\operatorname{\mathbb{S}}^{K}.$ This means that the RL algorithm may place limit orders up to K1=5K-1=5 ticks above the best bid price and observes the volumes on the first five levels. To encode the policy $\pi_{\theta}$ and the value function $V_{\vartheta},$ we use two separate feed-forward neural networks with two hidden layers, $\tanh$-activation functions in the hidden layers, and no activation functions in the last layers. The output layer of the policy network, specifying the mean $\mu_{\theta},$ has $K$ nodes, and the last layer of the value function network has one node. We initialize the covariance matrix $\Sigma$ to the identity matrix and then scale it down following the schedule (13). We initialize the bias of the output layer of the policy network to $b=(-1,-1,\dots,-1)\in\operatorname{\mathbb{R}}^{K}$ and set the other weights of the output layer of the policy network close to zero. Then, we obtain from (11) that
@@
$\operatorname{\mathbb{E}}\left[\log\left(\frac{a^{K}}{a^{k}}\right)\right]\approx 1\quad\text{and}\quad\operatorname{\mathbb{E}}\left[\log\left(\frac{a^{k}}{a^{K}}\right)\right]\approx-1,\quad k=0,1,\dots,K-1,$ (18)
@@
showing that action $a_{K}$ is more likely than actions $a_{0},\dots,a_{K-1}.$ Therefore, the algorithm is less likely to place orders at the beginning of the training cycle. This increases the likelihood of observing complete trajectories up to the terminal time $T$ early in the training process, thereby helping to prevent convergence to local minima. When we evaluate the algorithm after training is completed, we use a deterministic version of it. More precisely, we use the action $h(\mu_{\theta}),$ where $\mu_{\theta}$ is the mean of the underlying normal distribution, parametrized by the trained neural network weights, and $h$ is the logistic-normal transform; see Section 4.2. Further implementation details concerning the training of the LN algorithm are discussed in Appendix B.1.
We compare the LN algorithm with heuristic benchmarks and an algorithm based on the Dirichlet distribution (DR). For evaluation, we use the deterministic version of the DR algorithm, selecting the mean of the algorithm’s action distribution as the action. More implementation details of this algorithm are discussed in Appendix B.2.
In the following sections, we turn to a detailed discussion of the results. The test results are summarized in Table 1, which presents the expected value and standard deviation of the sum of normalized rewards defined in (3) across all market environments, position sizes, and execution algorithms. For example, the reward of 1.09-1.09 for the SL algorithm at 60 lots in the market with noise traders means that on average, the algorithm sold each of the 60 lots at 1.09-1.09 ticks less than the initial bid price. Figure 4 shows histograms of the sum of normalized rewards. Expected values, standard deviations, and reward histograms are calculated from 10,000 market simulations. The random seed for the market environments used to evaluate the algorithms differs from that used during training, thereby avoiding overfitting to the training simulations.
@@ 6.4 Logistic-normal vs Dirichlet distribution
We start by comparing the LN and DR algorithms. We observe that across all three market simulations, the LN algorithm outperforms the DR algorithm, except in the market with noise traders and in the market with noise and tactical traders when 60 lots are traded, as shown in Table 1. However, when the LN algorithm performs worse than the DR algorithm, the difference is only marginal. Figure 5 shows the average rewards per batch over the full training cycle. The upper panel shows the rewards for 20 lots, and the lower panel shows the rewards for 60 lots. We observe that the rewards for the LN algorithm converge better than for the DR algorithm, except when it sells 60 lots in the market that only contains noise traders. Overall, this suggests that the LN algorithm is a superior alternative to the DR algorithm for the trade execution problem.
Table 1: The table shows the expected rewards and their standard deviations for the submit and leave algorithm (SL), the time-weighted average price algorithm (TWAP), the logistic-normal algorithm (LN), and the Dirichlet algorithm (DR) in three simulated markets and for initial positions of 20 and 60 lots. The best expected value per row is highlighted in bold.
Market #Lots $\mathbb{E}[\text{SL}]$ $\sigma[\text{SL}]$ $\mathbb{E}[\text{TWAP}]$ $\sigma[\text{TWAP}]$ $\mathbb{E}[\text{DR}]$ $\sigma[\text{DR}]$ $\mathbb{E}[\text{LN}]$ $\sigma[\text{LN}]$
Noise 20 0.52 1.19 -0.06 0.94 0.21 0.76 0.61 1.03
60 -1.09 1.34 -1.40 0.98 -0.71 1.0 -0.72 0.90
Noise & Tactical 20 0.10 1.43 0.48 0.68 0.73 0.65 0.81 0.64
60 -3.36 0.99 -0.96 0.95 -0.23 0.65 -0.25 0.67
Noise & Tactical 20 -1.64 2.95 -0.36 3.03 1.06 2.24 1.13 2.08
& Strategic 60 -2.51 3.67 -1.45 3.46 0.03 1.97 0.23 2.15
@@
Figure 4: The different panels show reward distributions of the submit and leave algorithm (SL), the time-weighted average price algorithm (TWAP), the logistic-normal algorithm (LN) and the Dirichlet algorithm (DR) for three simulated markets. The first row corresponds to 20 lots, and the second to 60 lots.
@@
Figure 5: Average episode return per batch during training for different environments, initial position sizes, and algorithms. The x-axis shows the number of gradient steps. The top row shows the average returns for 20 lots. The bottom row shows the average returns for 60 lots. The green line corresponds to the LN algorithm, and the red line corresponds to the DR algorithm.
@@ 6.5 Market with noise traders
We note that all algorithms across all markets exhibit high standard deviation. This is due to the highly stochastic environment, combined with the fact that minimizing the standard deviation is not part of the training criterion. Sending a market order at the beginning of the execution period would result in a strategy with almost zero standard deviation; however, the execution price would be worse than the initial best bid and, therefore, worse than most of our execution strategies. The first two rows of Table 1 and the reward histograms in Figure 4(a) demonstrate that the LN algorithm outperforms the heuristic benchmark strategies for 20 and 60 lots in the market with noise traders. It is better than the DR algorithm for 20 lots and slightly worse for 60 lots. The rewards per batch in Figure 5(a) show much better convergence for the LN algorithm compared to the DR algorithm when it trades 20 lots, and similar convergence when it trades 60 lots. Interestingly, the reward histograms of the LN and DR algorithms differ when the initial position is 20 lots, with the LN algorithm’s histogram exhibiting more pronounced peaks at one and two ticks above the best bid price. This indicates that the algorithms employ different order placement strategies. For 60 lots, the LN algorithm’s histogram exhibits less skew to the left compared with the heuristic benchmark strategies, suggesting that it crosses the spread when the price moves down. The TWAP algorithm performs the worst for both position sizes. Since the market does not react to volume imbalance, placing the entire position passively at the start works relatively well. On the other hand, spreading the order over the entire execution interval leads to a worse performance for two reasons. First, due to the shape of the order book (Figure 8 below), limit orders at the best ask are less likely to be filled when the market is moving up, and more likely to be filled when the market is moving down. Second, it is more likely that the algorithm does not sell its inventory by the end of the execution period and is forced to cross the spread at terminal time.